OpenStudy (anonymous):

Is this correct! I have to find the domain for 12y-15/25y^2-4. I came up with 25y^2-4=0, then (5y-2)(5y+2)=0, then 5y-2=0 , 5y+2=0,adding 2 and -2 to both side equal to 5y=2 and 5y=-2 dividing by 5 gives me y=2/5 and y=-2/5

OpenStudy (zzr0ck3r):

so your domain is \(D=\{x\in \mathbb{R} \ | \ x\ne \frac{1}{5},\frac{-1}{5}\}\)

OpenStudy (zzr0ck3r):

all real numbers x such that x does not equal 1/5 or negative 1/5.

OpenStudy (anonymous):

What does the D={x=R|x=1/5, 1/5 means

jimthompson5910 (jim_thompson5910):

zzr0ck3r, it's 2/5, not 1/5

jimthompson5910 (jim_thompson5910):

that notation basically means the domain D is the set of all numbers such that x can be any real number but it cannot be equal to -2/5 or 2/5

OpenStudy (anonymous):

Thanks a lot.